Stocks have a 12% expected return and 22% risk.

Bonds have a 7% expected return and 10% risk.

The expected return of a portfolio comprised of 70% stocks and 30% bonds is:

A. 18.4%

B.10.5%

C.Cant tell from the information provided. You need to the return correlation between stocks% and bond.

D.8.8%

C. 7.2%
Answer #1

Expected return stock E(rs) = 12%

Expected return bond E(rb) = 7%

Weight of stock Ws = 0.70

Weight of bond Wb = 0.30

Expected return of the portfolio is weighted average of its assets.

So, Expected return of portfolio = Ws*E(rs) + Wb*E(rb) = 0.7*12 + 0.3*7 = 10.5%

So option B is correct.